OpenStudy (anonymous):

What is the solution to the equation 4 ^-2x ≈ 3 ?

OpenStudy (paxpolaris):

Why do you say approximately 3? Any way take the log of both sides to solve it...

OpenStudy (anonymous):

whick is equal to?

OpenStudy (anonymous):

\[-2x \log (4)=\log(3)\] \[-2x=\frac{\log3}{\log4}\] Then divide both sides by -2 to solve for x. There is indeed an exact answer.

OpenStudy (anonymous):

(I take the bottom part away).

OpenStudy (anonymous):

?????????

OpenStudy (paxpolaris):

\[\Large x= - {\log(3) \over 2 \cdot \log(4)}\] http://lmgtfy.com/?q=-log(3)%2F(2*log(4))

OpenStudy (anonymous):

Meaning?